OpenStudy (anonymous):

anyone know how (3x+11)/(x^2 - x- 6) is simplified to 4/(x-3) - 1/(x+2)?? I get that the denominator is factored, but how is the numerator changed?

OpenStudy (anonymous):

(3x+11)/(x-3)(x+2) = (4x+8-x+3)/(x-3)(x+2) (4(x+2)-(x-3))/(x-3)(x+2) = 4/(x-3) - 1/(x+2)